A plant pot of mass 3 kg falls from a balcony and drops 10.5 m to the ground. It falls freely under gravity. Find the speed of the plant pot just before it hits the ground.

Thanks
1
Expert's answer
2020-09-21T08:32:20-0400

According to the law of the conservation of energy:

"mgh=\frac{mv^2}{2}"

"gh=\frac{v^2}{2}"

"v^2=2gh"

"v=\sqrt{2gh}"

"v=\sqrt{2*10*10.5}=\sqrt{210}=14.5(\frac{m}{s})"